Can someone please help me with this problem? I am having trouble with figuring it out. Thank you in advanced! The 10x objective has a numerical aperture of 0.25. What would be the smallest distance (in μm) between two objects that can be resolved when using that objective with light at 400 nm?
Answer #1

The smallest distance between two objects that can be resolved by a microscope is known as resolution of that microscope and the resolution of a microscope is given by the formula: Resolution (r) = λ/(2NA) where λ = wavelength of light and NA = numerical aperture. Here the λ is given as 400 nm (0.4 μm) and NA is given as 0.25. Thus,

r = 0.4/ 0.25 = 1.6 μm

Thus, the smallest distance (in μm) between two objects that can be resolved is 1.6 in μm.

Note: Here answer is required in μm therefore wavelength of light (400 nm) has been converted into μm.